OpenStudy (anonymous):

Solve, using the substitution method. 3x + 2y = 11 x - 2 = -4 A. (0,-2) B. (4,-1/2) C. (12,-5/2) D. (0,-1/2)

OpenStudy (anonymous):

@satellite73 @hartnn

OpenStudy (anonymous):

@confluxepic @blurbendy @Joel_the_boss

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@just_one_last_goodbye @sleepyjess @sangya21 @I_Need_Help_With_Home @chucky00

OpenStudy (anonymous):

@hayhay5270 @Reaper534 @satellite73

OpenStudy (anonymous):

3x + 2y = 11 x - 2 = -4 x-2 = -4 x = -4+2 x = -2 put x in first eq. to find y

OpenStudy (secret-ninja):

http://www.webmath.com/solver2.html this site might help

OpenStudy (anonymous):

ok thank you :3

OpenStudy (anonymous):

@pumkinspicelatte! i will help in a min im gonna finish my project : )

OpenStudy (anonymous):

whata think ????

OpenStudy (confluxepic):

I was tagged here. @pumkinspicelatte! @~ Do you need more help. @confluxepic is here.